# Floquet breathers in a modulated nonlinear lattice

### v2

@@ B.2 Dispersion bands modification with only space modulation
In order to understand the result we will consider first the case $h=2\pi/L$, where $L$ is an integer, and $\Omega=0$. In this case, the symmetry of the system $n\rightarrow n+1$ is broken and substituted by the symmetry $n\rightarrow n+L$. Let us suppose that $\delta=0$, but still the symmetry is broken, then, in order to be able to impose periodic boundary conditions, we need that $N=RL$ with $R$ and integer number. Now, by Bloch theorem Griffiths and Schroeter (2018), the solutions are of the form $\exp(\textrm{i}qn)F(n,t)$, with $F(n,t)$ a solution with the periodicity of the lattice, whose period is now $L$, that is the elements of the basis are given by $u_{l,r}=\exp(\textrm{i}q_{r}n)\exp[\textrm{i}(q_{l}n-\omega t])$. As the second factor has the periodicity of the new lattice $q_{l}=(2\pi/L)l$, with $l=0,\dots,L-1$. If we impose periodic boundary conditions, then $u_{l,r}(n,t)=u_{l,r}(n+N,t)$, the second factor is automatically identical if $N=RL$ and the first factor becomes $q_{r}=2\pi r/N$, with $r=0,\dots,R-1$, as larger values of $r$ can be written as $q_{r}^{\prime}=2\pi(r_{1}R+r_{2})/N=2\pi(r_{1}R+r_{2})/LR=2\pi r_{1}/L+2\pi r_{2}/N$, and the first term is periodic in $L$ and can be incorporated into $F(n,t)$. We can rewrite it as:
@@
In the following, we use the standard notation of ket and bras summarized in Appendix A.
@@
[…]
@@
Figure 15: (Top:) Numerical dispersion bands for $h=2\pi/3$; (Bottom:) Theoretical degeneracy raising of the dispersion bands due to space modulation. Note that the Brillouin zone is $0,2\pi/3]$. See the text for explanation. Parameters: $\kappa=0.10.055794$, $\delta=0.05$, $\Omega=0$ .
B.3 Effect of only time modulation on the phonon bands
When $h=0$, the effect of time modulation is to split the phonon band in several parallel bands separated by a frequency of $\Omega$. They are obtained numerically as explained in App. B.1. Some bands are weak, meaning that their intensity in the XTFFT is small, presumably because their density of states is small. Some becomes so weak that they can not be seen into the background noise. The number of bands diminishes with increased frequency and eventually the bands disappear or are not visible. Extra bands appear for frequency below $\omega_{0}=1$ although they are very weak when approaching $\Omega=\omega_{0}$. Frequencies multiples of $\omega_{0}$ lead to a divergence of the system. Frequencies inside the phonon band excites the modes with that frequency.
These bands can be obtained analytically using the breakup of the time-invariance brought about by the time modulation. For that, we will use the notation and theory treated in Appendix A, taking into account the changes brought about by the time dependence of the on-site potential.
@@ B.3 Effect of only time modulation on the phonon bands
The new solution is given by Bloch theorem Slater (1958); Cassedy and Oliner (1963), as $\exp(-\textrm{i}\omega t)$, multiplied by a periodic function with the period of the operator, that is
@@
$(\omega+m\Omega)^{2}A_{m}=\bar{\omega}_{q}^{2}A_{m}+\frac{\displaystyle\delta}{\displaystyle 2}(A_{m+1}+A_{m-1})\,$, ($4$)
where $\bar{\omega}_{q}$ is the unperturbed frequency for the wavenumber $q$. This is a complicated equation, but if $\delta$ is small with respect to $\Omega$, we can suppress that term in the first approximation, that is.
@@
B.4 Dispersion bands for space-time modulation
Considering the system (7):
@@ B.4 Dispersion bands for space-time modulation
The system is now $2\pi$ periodic in the variable $hn-\Omega t$, and the solutions can be written as the product of a plane wave multiplied by a function with the periodicity of the system. That is:
$u=\exp(\textrm{i}[qn-\omega t])\sum_{m=-M/2}^{M/2}A_{m}\exp(\textrm{i}m[hn-\Omega t])\,.$
Then, the substitution of $u$ in the Hamiltonian leads to (substituting $\cos(hn-\Omega t)=(\exp(\textrm{i}[hn-\Omega t])+\exp(-\textrm{i}[hn-\Omega t))/2$:
@@
Assuming that $\delta$ is small with respect to $\Omega$, we can obtain the independent solutions where all the $A_{k}=0$, except one $A_{m}$ each time, so it corresponds to a single harmonic with $\omega$:

### v3

@@ C.2 Dispersion bands modification with only space modulation
In order to understand the result we will consider first the case $h=2\pi/L$, where $L$ is an integer, and $\Omega=0$. In this case, the symmetry of the system $n\rightarrow n+1$ is broken and substituted by the symmetry $n\rightarrow n+L$. Let us suppose that $\delta=0$, but still the symmetry is broken, then, in order to be able to impose periodic boundary conditions, we need that $N=RL$ with $R$ and integer number. Now, by Bloch theorem,Griffiths and Schroeter (2018) the solutions are of the form $\exp(\textrm{i}qn)F(n,t)$, with $F(n,t)$ a solution with the periodicity of the lattice, whose period is now $L$, that is the elements of the basis are given by $u_{l,r}=\exp(\textrm{i}q_{r}n)\exp[\textrm{i}(q_{l}n-\omega t])$. As the second factor has the periodicity of the new lattice $q_{l}=(2\pi/L)l$, with $l=0,\dots,L-1$. If we impose periodic boundary conditions, then $u_{l,r}(n,t)=u_{l,r}(n+N,t)$, the second factor is automatically identical if $N=RL$ and the first factor becomes $q_{r}=2\pi r/N$, with $r=0,\dots,R-1$, as larger values of $r$ can be written as $q_{r}^{\prime}=2\pi(r_{1}R+r_{2})/N=2\pi(r_{1}R+r_{2})/LR=2\pi r_{1}/L+2\pi r_{2}/N$, and the first term is periodic in $L$ and can be incorporated into $F(n,t)$. We can rewrite it as:
@@
In the following, we use the standard notation of ket and bras summarized in Appendix B.
@@
will be in the first approximation close to the eigenspace of $E_{0}$, that is some linear combination of $|k_{1}\rangle\equiv|k\rangle$ and $|k_{2}\rangle\equiv|-k\rangle$ (it can be done with the cosine and sine solutions but the theory has to be reformulated and does not add any value). So, let us suppose that $|\phi_{1}\rangle$ is a a linear combination of the orthogonal eigenvectors, i.e., $|\phi_{1}\rangle=\alpha|k_{1}\rangle+\beta|k_{2}\rangle$ and then, we obtain
@@
Figure 17: (Top:) Numerical dispersion bands for $h=2\pi/3$; (Bottom:) Theoretical degeneracy raising of the dispersion bands due to space modulation. Note that the Brillouin zone is $0,2\pi/3]$. See the text for explanation. Parameters: $\kappa=0.10$, $\delta=0.05$, $\Omega=0$ .
C.3 Effect of only time modulation on the phonon bands
When $h=0$, the effect of time modulation is to split the phonon band in several parallel bands separated by a frequency of $\Omega$. They are obtained numerically as explained in App. C.1. Some bands are weak, meaning that their intensity in the XTDFT is small, presumably because their density of states is small. Some becomes so weak that they can not be seen into the background noise. The number of bands diminishes with increased frequency and eventually the bands disappear or are not visible. Extra bands appear for frequency below $\omega_{0}=1$ although they are very weak when approaching $\Omega=\omega_{0}$. Frequencies multiples of $\omega_{0}$ lead to a divergence of the system. Frequencies inside the phonon band excites the modes with that frequency.
These bands can be obtained analytically using the breakup of the time-invariance brought about by the time modulation. For that, we will use the notation and theory treated in Appendix B, taking into account the changes brought about by the time dependence of the on-site potential.
@@ C.3 Effect of only time modulation on the phonon bands
The new solution is given by Bloch theorem,Slater (1958); Cassedy and Oliner (1963) as $\exp(-\textrm{i}\omega t)$, multiplied by a periodic function with the period of the operator, that is
@@
$(\omega+m\Omega)^{2}A_{m}=\bar{\omega}_{q}^{2}A_{m}+\frac{\displaystyle\delta}{\displaystyle 2}(A_{m+1}+A_{m-1})\,$, (51)
where $\bar{\omega}_{q}$ is the unperturbed frequency for the wavenumber $q$. This is a complicated equation, but if $\delta$ is small with respect to $\omega_{0}`2$, we can suppress that term in the first approximation, that is.
@@
C.4 Dispersion bands for space-time modulation
Considering the system (7) with both $\Omega\neq 0$ and $h\neq 0$:
@@ C.4 Dispersion bands for space-time modulation
The system is now $2\pi$ periodic in the variable $hn-\Omega t$, and the solutions can be written as the product of a plane wave multiplied by a function with the periodicity of the system.Cassedy and Oliner (1963) That is:
$u_{n}(t)=\exp(\textrm{i}[qn-\omega t])\sum_{m=-K/2}^{{K}/2}A_{m}\exp(\textrm{i}m[hn-\Omega t])\,$, ($5$)
with $K$ being an integer large enough to obtain the convergence of the truncated Fourier series. Then, the substitution of $u_{n}(t)$ in the Hamiltonian leads to (substituting $\cos(hn-\Omega t)=(\exp(\textrm{i}[hn-\Omega t])+\exp(-\textrm{i}[hn-\Omega t))/2$):
@@
Assuming that $\delta$ is small with respect to $\omega_{0}^{2}$, we can obtain the independent solutions where all the $A_{k}=0$, except one $A_{m}$ each time, so it corresponds to a single harmonic with $\omega$: